the work fuction of Na metal is 2.3 ev then the three shold wavelength lie in the which region?

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
As E = hc/ lambda
 
lambda = hc/E = 6.628 *10^-34 * 3 * 10^8 / 2.3 *1.6 * 10^-19
 
 = 5.403* 10^-7 m = 5403 Angstr